### v2

# C-Pack: Packaged Resources To Advance General Chinese Embedding

Abstract
We introduce C-Pack, a package of resources that significantly advance the field of general Chinese embeddings. C-Pack includes three critical resources. 1) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets. 2) C-MTP is a massive text embedding dataset curated from labeled and unlabeled Chinese corpora for training embedding models. 3) C-TEM is a family of embedding models covering multiple sizes. Our models outperform all prior Chinese text embeddings on C-MTEB by up to +10% upon the time of the release. We also integrate and optimize the entire suite of training methods for C-TEM. Along with our resources on general Chinese embedding, we release our data and models for English text embeddings. The English models achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data. All these resources are made publicly available at https://github.com/FlagOpen/FlagEmbedding.
1 Introduction
Text embedding is a long-standing topic in natural language processing and information retrieval. By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling Karpukhin et al. (2020); Lewis et al. (2020); Guu et al. (2020). The recent popularity of large language models (LLMs) has made text embeddings even more important. Due to the inherent limitations of LLMs, such as world knowledge and action space, external support via knowledge bases or tool use is necessary. Text embeddings are critical to connect LLMs with these external modules Borgeaud et al. (2022); Qin et al. (2023).
Figure 1: The C-Pack resources to support general Chinese embedding.
The wide variety of application scenarios calls for a single unified embedding model that can handle all kinds of usages (like retrieval, ranking, classification) in any application scenarios (e.g., question answering, language modeling, conversation). However, learning general-purpose text embeddings is much more challenging than task-specific ones. The following factors are critical:
$\bullet$ Data. The development of general-purpose text embeddings puts forward much higher demands on the training data in terms of scale, diversity, and quality. To achieve high discriminative power for the embeddings, it may take more than hundreds of millions of training instances Izacard et al. (2021); Ni et al. (2021b); Wang et al. (2022b), which is orders of magnitude greater than typical task-specific datasets, like MS MARCO Nguyen et al. (2016) and NLI Bowman et al. (2015); Williams et al. (2017). Besides scale, the training data needs to be collected from a wide range of sources so as to improve the generality across different tasks Izacard et al. (2021); Wang et al. (2022b). Finally, the augmentation of scale and diversity will probably introduce noise. Thus, the collected data must be properly cleaned before being utilized for the training of embeddings Wang et al. (2022b).
$\bullet$ Training. The training of general-purpose text embeddings depends on two critical elements: a well-suited backbone encoder and an appropriate training recipe. While one can resort to generic pre-trained models like BERT Devlin et al. (2018) and T5 (Raffel et al., 2020), the quality of text embedding can be substantially improved by pre-training with large-scale unlabeled data Izacard et al. (2021); Wang et al. (2022b). Further, instead of relying on a single algorithm, it takes a compound recipe to train general-purpose text embedding. Particularly, it needs embedding-oriented pre-training to prepare the text encoder Gao and Callan (2021), contrastive learning with sophisticated negative sampling to improve the embedding’s discriminability Qu et al. (2020), and instruction-based fine-tuning Su et al. (2022); Asai et al. (2022) to integrate different representation capabilities of text embedding.
$\bullet$ Benchmark. Another pre-requisite condition is the establishment of proper benchmarks, where all needed capabilities of text embeddings can be comprehensively evaluated. BEIR Thakur et al. (2021) provides a collection of 18 to evaluate the embedding’s general performances on different retrieval tasks, e.g., question answering and fact-checking. Later, MTEB Muennighoff et al. (2022a) proposes a more holistic evaluation of embeddings and extends BEIR. It integrates 56 datasets, where all important capabilities of text embeddings, like retrieval, ranking, clustering, etc., can be jointly evaluated.
Altogether, the development of general-purpose text embedding needs to be made on top of a mixture of driving forces, from data, and encoder models, to training methods and benchmarking. In recent years, continual progress has been achieved in this field, such as work from Contriever Izacard et al. (2021), E5 Wang et al. (2022b), and OpenAI Text Embedding Neelakantan et al. (2022). However, most of these works are oriented to the English world. In contrast, there is a severe shortage of competitive models for general Chinese embedding due to a series of limitations: there are neither well-prepared training resources nor suitable benchmarks to evaluate the generality.
To address the above challenges, we present a package of resources called C-Pack, which contributes to the development of general Chinese embedding from the following perspectives.
$\bullet$ C-MTEB (Chinese Massive Text Embedding Benchmark). The benchmark is established as a Chinese extension of MTEB. [^1] https://huggingface.co/spaces/mteb/leaderboard C-MTEB collects 35 public-available datasets belonging to 6 types of tasks. Thanks to the scale and diversity of C-MTEB, all major capabilities of Chinese embeddings can be reliably measured, making it the most suitable benchmark to evaluate the generality of Chinese text embedding.
$\bullet$ C-MTP (Chinese Massive Text Pairs). We create a massive training dataset of 100M text pairs, which integrates both labeled data and unlabeled data curated from Wudao Yuan et al. (2021), one of the largest corpora for pre-training Chinese language models. C-MTP is not only large and diverse but also cleaned to ensure the data quality.
$\bullet$ C-TEM (Chinese Text Embedding Models). We provide a family of well-trained models for Chinese general text embeddings. There are three optional model sizes: small (24M), base (102M), and large (326M), which present users with the flexibility to trade off efficiency and effectiveness. Our models make a big leap forward in generality: C-TEM outperforms all previously Chinese text embedding models on all aspects of C-MTEB by large margins. Besides being directly applicable, C-TEM can also be fine-tuned with additional data for better task-specific performances.
$\bullet$ Training Recipe. Accompanying our resources, we integrate and optimize training methods to build general-purpose text embeddings, including the pre-training of an embedding-oriented text encoder, general-purpose contrastive learning, and task-specific fine-tuning. The release of the training recipe will help the community to reproduce the state-of-the-art methods and make continuous progress on top of them.
In summary, C-Pack provides a go-to option for people’s application of general-purpose Chinese text embedding. It substantially advances the training and evaluation, laying a solid foundation for the future development of this field.
Figure 2: Overview of C-Pack. C-MTEB is a benchmark for Chinese text embeddings. C-MTP is a large-scale Chinese embedding training dataset. C-TEM are state-of-the-art Chinese embedding models. The training recipe is shown at the bottom.
2 C-Pack
In this section, we first introduce the resources in C-Pack: the benchmark C-MTEB, the training data C-MTP, and the model class C-TEM. Then, we discuss the training recipe, which enables us to train the state-of-the-art models for general Chinese embedding based on the offered resources.
2.1 Benchmark: C-MTEB
[…]
In particular, we collect a total of 35 public datasets, all of which can be used to evaluate Chinese text embeddings. The collected datasets are categorized based on the embedding’s capability they may evaluate. There are 6 groups of evaluation tasks: retrieval, re-ranking, STS (semantic textual similarity), classification, pair classification, and clustering, which cover the main interesting aspects of Chinese text embeddings. Note that there are multiple datasets for each category. The datasets of the same category are collected from different domains and complementary to each other, therefore ensuring the corresponding capability to be fully evaluated.
The nature of each task and its evaluation metric are briefly introduced as follows.
$\bullet$ Retrieval. The retrieval task is presented with the test queries and a large corpus. For each query, it finds the Top-$k$ similar documents within the corpus. The retrieval quality can be measured by ranking and recall metrics at different cut-offs. In this work, we use the setting from BEIR Thakur et al. (2021), using NDCG@10 as the main metric.
$\bullet$ Re-ranking. The re-ranking task is presented with test queries and their lists of candidate documents (one positive plus N negative documents). For each query, it re-ranks the candidate documents based on the embedding similarity. The MAP score is used as the main metric.
$\bullet$ STS (Semantic Textual Similarity). The STS (Agirre et al., 2012, 2013, 2014, 2015, 2016) task is to measure the correlation of two sentences based on their embedding similarity. Following the original setting in Sentence-BERT Reimers and Gurevych (2019), the Spearman’s correlation is computed with the given label, whose result is used as the main metric.
$\bullet$ Classification. The classification task re-uses the logistic regression classifier from MTEB Muennighoff et al. (2022a), where the provided label is predicted based on the input embedding. The average precision is used as the main metric.
$\bullet$ Pair-classification. This task deals with a pair of input sentences, whose relationship is presented by a binarized label. The relationship is predicted by embedding similarity, where the average precision is used as the main metric.
$\bullet$ Clustering. The clustering task is to group sentences into meaningful clusters. Following the original setting in MTEB Muennighoff et al. (2022a), it uses the mini-batch k-means method for the evaluation, with batch size equal to 32 and k equal to the number of labels within the mini-batch. The V-measure score is used as the main metric.
Finally, the embedding’s capability on each task is measured by the average performance of all datasets for that task. The embedding’s overall generality is measured by the average performance of all datasets in C-MTEB.
dataset
C-MTP (unlabeled)
C-MTP      (labeled)
source
Wudao, CSL, XLSUM-Zh, Amazon-Review-Zh, CMRC, etc.
[…]
size
[…]
Table 1: Composition of C-MTP
2.2 Training Data: C-MTP
We curate the largest dataset C-MTP for the training of general Chinese embedding. The paired texts constitute the data foundation for the training of text embedding, e.g., a question and its answer, two paraphrase sentences, or two documents on the same topic. To ensure the generality of the text embedding, the paired texts need to be both large-scale and diversified. Therefore, C-MTP is collected from two sources: the curation of massive unlabeled data, a.k.a. C-MTP (unlabeled); and the comprehensive collection of labeled data, a.k.a. C-MTP (labeled). The data collection process is briefly introduced as follows.
$\bullet$ C-MTP (unlabeled). We look for a wide variety of corpora, where we can extract rich-semantic paired structures from the plain text, e.g., paraphrases, title-body. Our primary source of data comes from open web corpora. The most representative one is the Wudao corpus Yuan et al. (2021), which is the largest well-formatted dataset for pre-training Chinese language models. For each of its articles, we extract (title, passage) to form a text pair. Following the same recipe, we also collect such text pairs from other similar web content like Zhihu, Baike, news websites, etc. Aside from the open web content, we also explore other public Chinese datasets to extract text pairs, such as CSL (scientific literature), Amazon-Review-Zh (reviews), Wiki Atomic Edits (paraphrases), CMRC (machine reading comprehension), XLSUM-Zh (summarization), etc. The paired structures are obvious in these datasets, which are directly extracted for the augmentation of C-MTP (unlabeled).
The text pairs curated from the web and other public sources are not guaranteed to be closely related. Therefore, data quality can be a major concern. In our work, we use a simple strategy to filter the data before adding it to C-MTP (unlabeled). Particularly, we use a third-party model: Text2Vec-Chinese [^2] https://huggingface.co/GanymedeNil to score the strength of relation for each text pair. We empirically choose a threshold of 0.43, and drop the samples whose scores are below the threshold. With such an operation, there are 100 million text pairs filtered from the unlabeled corpora. Despite the simplicity, we find that it effectively removes the irrelevant text pairs when manually reviewing samples and leads to strong empirical performances for the models trained on C-MTP (unlabeled).
[…]
Given the differences in scale and quality, C-MTP (unlabeled) and C-MTP (labeled) are applied to different training stages, which jointly result in a strong performance for the embedding model. Detailed analysis will be made in our training recipe.
2.3 Model Class: C-TEM
We provide a comprehensive class of well-trained embedding models for the community. Our models take a BERT-like architecture, where the last layer’s hidden state of the special token [CLS] is trained to work as the embedding. There are three different scales for the models: large (with 326M parameters), base (with 102M parameters), and small (with 24M parameters). The large-scale model achieves the highest general representation performances, leading the current public-available models by a considerable margin. The small-scale model is also empirically competitive compared with the public-available models and other model options in C-TEM; besides, it is way faster and lighter, making it suitable to handle massive knowledge bases and high-throughput applications. Thanks to the comprehensive coverage of different model sizes, people are presented with the flexibility to trade off running efficiency and representation quality based on their own needs.
As introduced, the models within C-TEM have been well-trained and achieve a strong generality for a wide variety of tasks. Meanwhile, they can also be further fine-tuned if 1) the embeddings are applied for a specific scenario, 2) the training data is presented for the application scenario. It is empirically verified that the fine-tuned model may bring forth a much better performance for its application, compared with its original model in C-TEM, and the fine-tuned models from other general pre-trained encoders, like BERT. In other words, C-TEM not only presents people with direct usage embeddings but also works as a foundation where people may develop more powerful embeddings.
2.4 Training Recipe
The training recipe of C-TEM is completely released to the public along with C-Pack (Figure 2). Our training recipe has three main components: 1) pre-training with plain texts, 2) contrastive learning with C-MTP (unlabeled), and 3) multi-task learning with C-MTP (labeled), whose specifications are made as follows.
$\bullet$ Pre-Training. Our model is pre-trained on massive plain texts through a tailored algorithm in order to better support the embedding task. Particularly, we make use of the Wudao corpora Yuan et al. (2021), which is a huge and high-quality dataset for Chinese language model pre-training. We leverage the MAE-style approach presented in RetroMAE Liu and Shao (2022); Xiao et al. (2023), which is simple but highly effective. The polluted text is encoded into its embedding, from which the clean text is recovered on top of a light-weight decoder:
$\min.\sum_{x\in X}-\log\mathrm{Dec}(x|\mathbf{e}_{\tilde{X}}),~{}\mathbf{e}_{\tilde{X}}\leftarrow\mathrm{Enc}(\tilde{X}).$
($\mathrm{Enc}$, $\mathrm{Dec}$ are the encoder and decoder, $X$, $\tilde{X}$ indicate the clean and polluted text.)
$\bullet$ General purpose fine-tuning. The pre-trained model is fine-tuned on C-MTP (unlabeled) via contrastive learning, where it is learned to discriminate the paired texts from their negative samples:
$\min.\sum_{(p,q)}-\log\frac{e^{\langle\mathbf{e}_{p},\mathbf{e}_{q}\rangle/\tau}}{e^{\langle\mathbf{e}_{p},\mathbf{e}_{q}\rangle/\tau}+\sum_{Q^{\prime}}e^{\langle\mathbf{e}_{p},\mathbf{e}_{q^{\prime}}\rangle/\tau}}.$
($p$ and $q$ are the paired texts, $q^{\prime}\in Q^{\prime}$ is a negative sample, $\tau$ is the temperature). One critical factor of contrastive learning is the negative samples. Instead of mining hard negative samples on purpose, we purely rely on in-batch negative samples Karpukhin et al. (2020) and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
model
Dim
Retrieval
STS
Pair CLF
CLF
Re-rank
Cluster
Average
Text2Vec (base)
768
38.79
43.41
67.41
62.19
49.45
37.66
48.59
Text2Vec (large)
1024
41.94
44.97
70.86
60.66
49.16
30.02
48.56
Luotuo (large)
1024
44.40
42.79
66.62
61.0
49.25
44.39
50.12
M3E (base)
768
56.91
50.47
63.99
67.52
59.34
47.68
57.79
M3E (large)
1024
54.75
50.42
64.30
68.20
59.66
48.88
57.66
Multi. E5 (base)
768
61.63
46.49
67.07
65.35
54.35
40.68
56.21
Multi. E5 (large)
1024
63.66
48.44
69.89
67.34
56.00
48.23
58.84
OpenAI-Ada-002
1536
52.00
43.35
69.56
64.31
54.28
45.68
53.02
BGE (small)
512
63.07
49.45
70.35
63.64
61.48
45.09
58.28
BGE (base)
768
69.53
54.12
77.50
67.07
64.91
47.63
62.80
BGE (large)
1024
71.53 54.98 78.94 68.32 65.11
48.39
63.96
Table 2: Performance of various models on C-MTEB.
$\bullet$ Task-specific fine-tuning. The embedding model is further fine-tuned with C-MTP (labeled). The labeled datasets are smaller but of higher quality. However, the contained tasks are of different types, whose impacts can be mutually contradicted. In this place, we apply two strategies to mitigate this problem. On one hand, we leverage instruction-based fine-tuning Su et al. (2022); Asai et al. (2022), where the input is differentiated to help the model accommodate different tasks. For each text pair ($p$, $q$), a task specific instruction $I_{t}$ is attached to the query side: $q^{\prime}\leftarrow q+I_{t}$. The instruction is a verbal prompt, which specifies the nature of the task, e.g., “search relevant passages for the query”. On the other hand, the negative sampling is updated: in addition to the in-batch negative samples, one hard negative sample $q^{\prime}$ is mined for each text pair ($p$, $q$). The hard negative sample is mined from the task’s original corpus, following the ANN-style sampling strategy in Xiong et al. (2020).
3 Experiments
We conduct experimental studies for the following purposes. P1. The extensive evaluation of different Chinese text embeddings on C-MTEB. P2. The empirical verification of the text embeddings by C-TEM. P3. The exploration of the practical value brought by C-MTP. P4. The exploration of the impacts introduced by the training recipe.
We consider the following popular Chinese text embedding models as the baselines for our experiments: Text2Vec-Chinese [^4] https://huggingface.co/shibing624 base and large; Luotuo [^5] https://huggingface.co/silk-road/luotuo-bert-medium; M3E [^6] https://huggingface.co/moka-ai base and large; multilingual E5 Wang et al. (2022b) and OpenAI text embedding ada 002777https://platform.openai.com/docs/guides/embeddings. The main metric presented in Section 2.1 is reported for each task in C-MTEB.
3.1 General Evaluation
We extensively evaluate C-TEM against popular Chinese text embeddings on C-MTEB as shown in Table 2.888Our C-TEM models are named BGE in the tables. We make the following observations.
First, our models outperform existing Chinese text embeddings by large margins. There is not only an overwhelming advantage in terms of the average performance, but also notable improvements for the majority of tasks in C-MTEB. The biggest improvements are on the retrieval task followed by STS, pair classification, and re-ranking. Such aspects are the most common functionalities of text embeddings, which are intensively utilized in applications like search engines, open-domain question answering, and the retrieval augmentation of large language models. Although the advantages for classification and clustering tasks are not as obvious, our performances are still on par or slightly better than the other most competitive models. The above observations verify the strong generality of C-TEM. Our models can be directly utilized to support different types of application scenarios.
Second, we observe performance growth resulting from the scaling up model size and embedding dimension. Particularly, the average performance improves from 58.28 to 63.96, when the embedding model is expanded from small to large. Besides the growth in average performance, there are also improvements across all the evaluation tasks. Compared to the other two baselines (Text2Vec, M3E), the impact of scaling up is more consistent and significant for our models. It is worth noting that our small model is still empirically competitive despite its highly reduced model size, where the average performance is even higher than the large-scale option of many existing models. As a result, it provides people with the flexibility to trade-off embedding quality and running efficiency: people may resort to our large-scale embedding model to deal with high-precision usages, or switch to the small-scale one for high-throughput scenarios.
model
Dim
Retrieval
STS
Pair CLF
CLF
Re-rank
Cluster
Average
M3E (large)
1024
54.75
50.42
64.30
68.20
59.66
48.88
57.66
OpenAI-Ada-002
1536
52.00
43.35
69.56
64.31
54.28
45.68
53.02
w.o. Instruct
1024
70.55
53.00
76.77
68.58
64.91
50.01
63.40
BGE-$i$
1024
63.90
47.71
61.67
68.59
60.12
47.73
59.00
BGE-$i$ w.o. pre-train
1024
62.56
48.06
61.66
67.89
61.25
46.82
58.62
BGE-$f$
1024
71.53 54.98 78.94
68.32
65.11
48.39
63.96
Table 3: Ablation of the training data, C-MTP, and the training recipe.
3.2 Detailed Analysis
We investigate the detailed impact of C-MTP and our training recipe. The corresponding experiment results are presented in Table 3 and Table 4.
First of all, we analyze the impact of our training data, C-MTP. As mentioned, C-MTP consists of two parts. 1) C-MTP (unlabeled), which is used for general-purpose fine-tuning; the model produced from this stage is called the intermediate checkpoint, denoted as BGE-$i$. 2) C-MTP (labeled), where the task-specific fine-tuning is further conducted on top of BGE-$i$; the model produced from this stage is called the final checkpoint, noted as BGE-$f$. Based on our observations from the experimental result, both C-MTP (unlabeled) and C-MTP (labeled) substantially contribute to the embedding’s quality.
Regarding C-MTP (unlabeled), despite mostly being curated from unlabeled corpora, this dataset alone brings forth strong empirical performance for the embedding models trained on it. Compared with other baselines like Text2Vec, M3E, and OpenAI text embedding, BGE-$i$ already achieves a higher average performance. A further look into the performances reveals more details. On one hand, C-MTP (unlabeled) makes a major impact on the embedding’s retrieval quality, where BGE-$i$ notably outperforms the baselines in this attribute. On the other hand, the general capability of embedding is primarily established with C-MTP (unlabeled), as BGE-$i$’s performance is close to the baselines on the rest of the aspects, like STS and Clustering. This puts our embedding models in a very favorable position for further improvements.
As for C-MTP (labeled), the dataset is much smaller but of better quality. With another round of fine-tuning on C-MTP (labeled), the empirical advantage is significantly expanded for the final checkpoint BGE-$f$, where it gives rise to a jump in average performance from 59.0 (BGE-$i$) to 63.96 (BGE-$f$). Knowing that the text pairs in C-MTP (labeled) are mainly gathered from retrieval and NLI tasks, the most notable improvements are achieved on closely related tasks, namely retrieval, re-ranking, STS, and pair classification. On other tasks, it preserves or marginally improves performance. This indicates that a mixture of high-quality and diversified labeled data is able to bring forth substantial and comprehensive improvements for a pre-trained embedding model.
We further explore the impact of our training recipe, particularly contrastive learning, task-specific fine-tuning, and pre-training.
One notable feature of our training recipe is that we adopt a large batch size for contrastive learning. According to previous studies, the learning of the embedding model may benefit from the increasing of negative samples Izacard et al. (2021); Qu et al. (2020); Muennighoff (2022). Given our dependency on in-batch negative samples, the batch size needs to be expanded as much as possible. In our implementation, we use a compound strategy of gradient checkpointing and cross-device embedding sharing Gao et al. (2021b), which results in a maximum batch size of 19,200. By making a parallel comparison between bz: 256, 2028, 19,200, we observe consistent improvement in embedding quality with the expansion of batch size (noted as bz). The most notable improvement is achieved in retrieval performance. This is likely due to the fact that retrieval is usually performed over a large database, where embeddings need to be highly discriminative.
Another feature is the utilization of instructions during task-specific fine-tuning. The task-specific instruction serves as a hard prompt. It differentiates the embedding model’s activation, which lets the model better accommodate a variety of different tasks. We perform the ablation study by removing this operation, noted as “w.o. Instruct”. Compared with this variation, the original method BGE-f gives rise to better average performance. Besides, there are more significant empirical advantages on retrieval, STS, pair classification, and re-rank. All these perspectives are closely related to the training data at the final stage, i.e. C-MTP (labeled), where the model is fine-tuned on a small group of tasks. This indicates that using instructions may substantially contribute to task-specific fine-tuning.
One more characteristic is that we use a specifically pre-trained text encoder to train C-TEM, rather than using common choices, like BERT and RoBERTa Liu et al. (2019). To explore its impact, we replace the pre-trained text encoder with the widely used Chinese-RoBERTa999huggingface.co/hfl/chinese-roberta-wwm-ext-large, noted as “BGE-$i$ w.o. pre-train”. According to the comparison with BGE-$i$, the pre-trained text encoder notably improves the retrieval capability, while preserving similar performances on other aspects.
Batch Size
256
2,048
19,200
Retrieval
57.25
60.96
63.90
STS
46.16
46.60
47.71
Pair CLF
62.02
61.91
61.67
CLF
65.71
67.42
68.59
Re-rank
58.59
59.98
60.12
Cluster
49.52
49.04
47.73
Average
56.43
57.92
59.00
Table 4: Impact of batch size.
4 Related Work
The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering Karpukhin et al. (2020) but also due to its fundamental role in augmenting large language models Lewis et al. (2020); Guu et al. (2020); Borgeaud et al. (2022); Izacard et al. (2022); Shi et al. (2023). Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios. In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever Izacard et al. (2021), GTR Ni et al. (2021b), sentence-T5 Ni et al. (2021a), Sentence-Transformer Reimers and Gurevych (2019), E5 Wang et al. (2022a), OpenAI text embedding Neelakantan et al. (2022), etc. Although it remains an open problem, recent studies highlight the following important factors. Firstly, the training data is desired to be large-scale and diversified, from which the embedding model can learn to recognize different kinds of semantic relationships Izacard et al. (2021); Wang et al. (2022b); Neelakantan et al. (2022). Secondly, the embedding model must be scaled up, as large text encoders are more generalizable across different application scenarios Muennighoff (2022); Ni et al. (2021b, a) in line with observations for the importance of scaling LLMs (Hoffmann et al., 2022; Rae et al., 2021; Brown et al., 2020; Chowdhery et al., 2022; Srivastava et al., 2022; Gao et al., 2021a; Li et al., 2023a; Allal et al., 2023; Muennighoff et al., 2023b). Thirdly, the training recipe must be optimized through pre-training Liu and Shao (2022); Wang et al. (2022a), negative sampling Izacard et al. (2021); Wang et al. (2022a), and multi-task fine-tuning Su et al. (2022); Asai et al. (2022); Sanh et al. (2021); Wei et al. (2021); Muennighoff et al. (2022b, 2023a); Chung et al. (2022). Aside from the above, it is also critical to establish proper benchmarks to evaluate the generality of text embeddings. Unlike previous task-specific evaluations, like MSMARCO Nguyen et al. (2016), SentEval Conneau and Kiela (2018), it is needed to substantially augment the benchmarks so as to evaluate the embedding’s performance for a wide variety of tasks. One representative work is made by BEIR Thakur et al. (2021); Kamalloo et al. (2023), where the embeddings can be evaluated across different retrieval tasks. It is later extended by MTEB Muennighoff et al. (2022a), where all major aspects of text embeddings can be comprehensively evaluated.
Given the above analysis, it can be concluded that the general text embedding is highly resource-dependent, which calls for a wide range of elements, such as datasets, models, and benchmarks. Thus, the creation and public release of the corresponding resources is crucially important.
5 Conclusion
We present C-Pack to advance progress towards general Chinese embedding. C-Pack consists of three core resources 1) The benchmark C-MTEB, covering 6 major tasks of embeddings and 35 datasets, making it the most comprehensive benchmark to evaluate the generality of Chinese embeddings. 2) The training data C-MTP, curated from massive unlabeled corpora and high-quality labeled datasets. Its unprecedented scale, diversity, and quality contribute to the superior generality of our embedding models. 3) The models C-TEM, which are empirically competitive. Their different sizes provide people with the flexibility to trade off efficiency and embedding quality. The entire training recipe is also provided along with these resources. The public release of C-Pack facilitates the usage of general Chinese embedding and also paves the way for its future advancement.
Appendix A C-MTEB Datasets
Table 5: Overview of datasets in C-MTEB.
Name URL Description Categ. Test        Samples
Classification
AmazonReviewsClassification Muennighoff et al. (2022a); McAuley and Leskovec (2013)
https://hf.co/datasets/mteb/amazon_reviews_multi
Sentiment of Amazon reviews
s2s
5,000
IFlyTek Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/IFlyTek-classification
Long text classification for App descriptions
s2s
2,600
JDReview (https://hf.co/datasets/kuroneko5943/jd21)
https://hf.co/datasets/C-MTEB/JDReview-classification
iPhone reviews
s2s
533
MassiveIntentClassification Muennighoff et al. (2022a); FitzGerald et al. (2022)
https://hf.co/datasets/mteb/amazon_massive_intent
Amazon Alexa virtual assistant utterances annotated with the associated intent
s2s
16,500
MassiveScenarioClassification Muennighoff et al. (2022a); FitzGerald et al. (2022)
https://hf.co/datasets/mteb/amazon_massive_scenario
Amazon Alexa virtual assistant utterances annotated with the associated scenario
s2s
16,500
MultilingualSentiment McAuley and Leskovec (2013)
https://hf.co/datasets/C-MTEB/MultilingualSentiment-classification
Sentiment of Amazon reviews
s2s
3,000
OnlineShopping (https://github.com/SophonPlus/ChineseNlpCorpus/blob/master/datasets/online_shopping_10_cats/intro.ipynb)
https://hf.co/datasets/C-MTEB/OnlineShopping-classification
Sentiment Analysis of User Reviews on Online Shopping Websites
s2s
1,000
TNews Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/TNews-classification
Short Text Classification for News
s2s
10,000
Waimai (https://github.com/SophonPlus/ChineseNlpCorpus/blob/master/datasets/waimai_10k/intro.ipynb)
https://hf.co/datasets/C-MTEB/waimai-classification
Sentiment Analysis of user reviews on takeaway platforms
s2s
1,000
Clustering
CLSClusteringP2P Li et al. (2022)
https://hf.co/datasets/C-MTEB/CLSClusteringP2P
Clustering of titles + abstract from CLS dataset. Clustering of 13 sets, based on the main category.
p2p
10,000
CLSClusteringS2S Li et al. (2022)
https://hf.co/datasets/C-MTEB/C-MTEB/CLSClusteringS2S
Clustering of titles from CLS dataset. Clustering of 13 sets, based on the main category.
s2s
10,000
ThuNewsClusteringP2P Li et al. (2006); Li and Sun (2007)
https://hf.co/datasets/C-MTEB/ThuNewsClusteringP2P
Clustering of titles + abstract from the THUCNews dataset
p2p
10,000
ThuNewsClusteringS2S Li et al. (2006); Li and Sun (2007)
https://hf.co/datasets/C-MTEB/ThuNewsClusteringS2S
Clustering of titles from the THUCNews dataset
s2s
10,000
Pair Classification
Cmnli Xu et al. (2020a, b); Conneau and Kiela (2018); Williams et al. (2017)
https://hf.co/datasets/C-MTEB/CMNLI
Chinese Multi-Genre NLI
s2s
139,000
Ocnli Hu et al. (2020)
https://hf.co/datasets/C-MTEB/OCNLI
Original Chinese Natural Language Inference dataset
s2s
3,000
Reranking
T2Reranking Xie et al. (2023)
https://hf.co/datasets/C-MTEB/T2Reranking
T2Ranking: A large-scale Chinese Benchmark for Passage Ranking
s2p
24,382
MMarcoRetrieval Bonifacio et al. (2021)
https://hf.co/datasets/C-MTEB/Mmarco-reranking
mMARCO is a multilingual version of the MS MARCO passage ranking dataset
s2p
7,437
CMedQAv1 Zhang et al. (2017)
https://hf.co/datasets/C-MTEB/CMedQAv1-reranking
Chinese community medical question answering
s2p
2,000
CMedQAv2 Zhang et al. (2018b)
https://hf.co/datasets/C-MTEB/C-MTEB/CMedQAv2-reranking
Chinese community medical question answering
s2p
4,000
Retrieval
T2Retrieval Xie et al. (2023)
https://hf.co/datasets/C-MTEB/T2Retrieval
T2Ranking: A large-scale Chinese Benchmark for Passage Ranking
s2p
24,832
MMarcoRetrieval Bonifacio et al. (2021)
https://hf.co/datasets/C-MTEB/MMarcoRetrieval
mMARCO is a multilingual version of the MS MARCO passage ranking dataset
s2p
7,437
DuRetrieval Qiu et al. (2022)
https://hf.co/datasets/C-MTEB/DuRetrieval
A Large-scale Chinese Benchmark for Passage Retrieval from Web Search Engine
s2p
4,000
CovidRetrieval Qiu et al. (2022)
https://hf.co/datasets/C-MTEB/CovidRetrieval
COVID-19 news articles
s2p
949
CmedqaRetrieval Qiu et al. (2022)
https://hf.co/datasets/C-MTEB/CmedqaRetrieval
Online medical consultation text
s2p
3,999
EcomRetrieval Long et al. (2022)
https://hf.co/datasets/C-MTEB/EcomRetrieval
Passage retrieval dataset collected from Alibaba search engine systems in e-commerce domain
s2p
1,000
MedicalRetrieval Long et al. (2022)
https://hf.co/datasets/C-MTEB/MedicalRetrieval
Passage retrieval dataset collected from Alibaba search engine systems in medical domain
s2p
1,000
VideoRetrieval Long et al. (2022)
https://hf.co/datasets/C-MTEB/VideoRetrieval
Passage retrieval dataset collected from Alibaba search engine systems in video domain
s2p
1,000
STS
AFQMC Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/AFQMC
Ant Financial Question Matching Corpus
s2s
3,861
ATEC (https://github.com/IceFlameWorm/NLP_Datasets/tree/master/ATEC)
https://hf.co/datasets/C-MTEB/ATEC
ATEC NLP sentence pair similarity competition
s2s
20,000
BQ Chen et al. (2018)
https://hf.co/datasets/C-MTEB/BQ
Bank Question Semantic Similarity
s2s
10,000
LCQMC Liu et al. (2018)
https://hf.co/datasets/C-MTEB/LCQMC
A large-scale Chinese question matching corpus.
s2s
12,500
PAWSX Yang et al. (2019)
https://hf.co/datasets/C-MTEB/PAWSX
Translated PAWS evaluation pairs
s2s
2,000
QBQTC Xu et al. (2020a)
https://hf.co/datasets/C-MTEB/QBQTC
QQ Browser Query Title Corpus
s2s
5,000
STSB Cer et al. (2017)
https://hf.co/datasets/C-MTEB/STSB
Translate STS-B into Chinese
s2s
1,360
STS-22 Muennighoff et al. (2022a)
https://hf.co/datasets/mteb/sts22-crosslingual-sts
Chinese news
p2p
656
Appendix B C-MTP Composition
We mine large-scale pairs of data from various domains. Table 6 shows the details for each data.
data source
type of text pairs
# of pairs
URL
cmrc2018
(query, context)
9,669
https://huggingface.co/datasets/cmrc2018
dureader
(query, context)
96,486
https://github.com/baidu/DuReader
simclue
[…]
388,779
https://github.com/CLUEbenchmark/SimCLUE
csl
(title, abstract)
394,846
https://arxiv.org/abs/2209.05034
amazon_reviews_multi
(title, body)
157,762
https://huggingface.co/datasets/amazon_reviews_multi
wiki_atomic_edits
(sentence, edited sentence)
1,213,688
https://huggingface.co/datasets/wiki_atomic_edits
mlqa
(question, context)
70,594
https://huggingface.co/datasets/mlqa
xlsum
(title, summary) (title, text)
89,505
https://huggingface.co/datasets/csebuetnlp/xlsum
wudao
(title, passage)
37,318,330
https://data.baai.ac.cn/details/WuDaoCorporaText
Misc
60,260,341
Table 6: Details for each dataset. The Misc data comes from the Internet, including QA, paper, and news data.
Appendix C English Models
Using our recipe, we also train a set of English text embedding models presented in Table 7. At the time of writing, our English BGE models are state-of-the-art on the English MTEB benchmark (Muennighoff et al., 2022a) across its 56 datasets. Our models outperform significantly larger models, such as SGPT Bloom which has 7.1 billion parameters (Muennighoff, 2022; Scao et al., 2022a, b). We advance the prior state-of-the-art by an absolute 1.1 (Li et al., 2023b). Our training recipe is the same as for our Chinese models, except for the usage of English data. We first finetune on unsupervised datasets including datasets like Wikipedia, CC-net, StackExchange, Reddit, S2orc, and datasets from sentence-transformers. [^10] https://huggingface.co/datasets/sentence-transformers/embedding-training-data We then further fine-tune on supervised datasets including NLI (Gao et al., 2021c), FEVER (Thorne et al., 2018), NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), Quora, StackExchange Duplicates and MEDI Su et al. (2022).
Model Name Dim. Average Retrieval Cluster Pair CLF Re-rank STS Summarize CLF
BGE (large) 1024 64.23 54.29 46.08 87.12 60.03 83.11 31.61 75.97
BGE (base) 768 63.55 53.25 45.77 86.55 58.86 82.4 31.07 75.53
BGE (small) 384 62.17 51.68 43.82 84.92 58.36 81.59 30.12 74.14
GTE (large) 1024 63.13 52.22 46.84 85.00 59.13 83.35 31.66 73.33
GTE (base) 768 62.39 51.14 46.2 84.57 58.61 82.3 31.17 73.01
E5 (large) 1024 62.25 50.56 44.49 86.03 56.61 82.05 30.19 75.24
Instructor-XL 768 61.79 49.26 44.74 86.62 57.29 83.06 32.32 61.79
E5 (base) 768 61.5 50.29 43.80 85.73 55.91 81.05 30.28 73.84
GTE (small) 384 61.36 49.46 44.89 83.54 57.7 82.07 30.42 72.31
OpenAI Ada 002 1536 60.99 49.25 45.9 84.89 56.32 80.97 30.8 70.93
E5 (small) 384 59.93 49.04 39.92 84.67 54.32 80.39 31.16 72.94
ST5 (XXL) 768 59.51 42.24 43.72 85.06 56.42 82.63 30.08 73.42
MPNet (base) 768 57.78 43.81 43.69 83.04 59.36 80.28 27.49 65.07
SGPT Bloom (7.1B) 4096 57.59 48.22 38.93 81.9 55.65 77.74 33.60 66.19
Table 7: Performance of English Models on MTEB.

### v3

@@
-C-Pack: Packaged Resources To Advance General Chinese Embedding
-Abstract
-We introduce C-Pack, a package of resources that significantly advance the field of general Chinese embeddings.
+C-Pack: Packed Resources For General Chinese Embeddings
+Abstract.
+We introduce C-Pack, a package of resources that significantly advance the field of general text embeddings for Chinese.
 C-Pack includes three critical resources.
-1) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets.
-2) C-MTP is a massive text embedding dataset curated from labeled and unlabeled Chinese corpora for training embedding models.
-3) C-TEM is a family of embedding models covering multiple sizes.
-Our models outperform all prior Chinese text embeddings on C-MTEB by up to +10% upon the time of the release.
-We also integrate and optimize the entire suite of training methods for C-TEM.
+1) C-MTP is a massive training dataset for text embedding, which is based on the curation of vast unlabeled corpora and the integration of high-quality labeled corpora.
+2) C-MTEB is a comprehensive benchmark for Chinese text embeddings covering 6 tasks and 35 datasets.
+3) BGE is a family of embedding models covering multiple sizes.
+Our models outperform all prior Chinese text embeddings on C-MTEB by more than +10% upon the time of the release.
+We also integrate and optimize the entire suite of training methods for BGE.
 Along with our resources on general Chinese embedding, we release our data and models for English text embeddings.
-The English models achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data.
+The English models also achieve state-of-the-art performance on the MTEB benchmark; meanwhile, our released English data is 2 times larger than the Chinese data.
+Both Chinese and English datasets are the largest public release of training data for text embeddings.
 All these resources are made publicly available at https://github.com/FlagOpen/FlagEmbedding.
-1 Introduction
+Text Embeddings, Training Data, Benchmark, Pre-trained Models
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Retrieval models and ranking
+1. Introduction
 Text embedding is a long-standing topic in natural language processing and information retrieval.
-By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling Karpukhin et al.
-(2020); Lewis et al.
-(2020); Guu et al.
-(2020).
+By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Xiong et al., 2020).
 The recent popularity of large language models (LLMs) has made text embeddings even more important.
 Due to the inherent limitations of LLMs, such as world knowledge and action space, external support via knowledge bases or tool use is necessary.
-Text embeddings are critical to connect LLMs with these external modules Borgeaud et al.
-(2022); Qin et al.
-(2023).
-Figure 1: The C-Pack resources to support general Chinese embedding.
+Text embeddings are critical to connect LLMs with these external modules (Borgeaud et al., 2022; Qin et al., 2023).
+Figure 1.
+C-Pack presents 4 critical resources to support general Chinese embedding: C-MTEB (comprehensive evaluation benchmark), C-MTP (massive training data), BGE (powerful pre-trained models), the entire-suite of training recipe.
 The wide variety of application scenarios calls for a single unified embedding model that can handle all kinds of usages (like retrieval, ranking, classification) in any application scenarios (e.g., question answering, language modeling, conversation).
@@
 The development of general-purpose text embeddings puts forward much higher demands on the training data in terms of scale, diversity, and quality.
-To achieve high discriminative power for the embeddings, it may take more than hundreds of millions of training instances Izacard et al.
-(2021); Ni et al.
-(2021b); Wang et al.
-(2022b), which is orders of magnitude greater than typical task-specific datasets, like MS MARCO Nguyen et al.
-(2016) and NLI Bowman et al.
-(2015); Williams et al.
-(2017).
-Besides scale, the training data needs to be collected from a wide range of sources so as to improve the generality across different tasks Izacard et al.
-(2021); Wang et al.
-(2022b).
+To achieve high discriminative power for the embeddings, it may take more than hundreds of millions of training instances (Izacard et al., 2021; Ni et al., 2021b; Wang et al., 2022b), which is orders of magnitude greater than typical task-specific datasets, like MS MARCO (Nguyen et al., 2016) and NLI (Bowman et al., 2015; Williams et al., 2017).
+Besides scale, the training data needs to be collected from a wide range of sources so as to improve the generality across different tasks (Izacard et al., 2021; Wang et al., 2022b).
 Finally, the augmentation of scale and diversity will probably introduce noise.
-Thus, the collected data must be properly cleaned before being utilized for the training of embeddings Wang et al.
-(2022b).
+Thus, the collected data must be properly cleaned before being utilized for the training of embeddings (Wang et al., 2022b).
 $\bullet$ Training.
 The training of general-purpose text embeddings depends on two critical elements: a well-suited backbone encoder and an appropriate training recipe.
-While one can resort to generic pre-trained models like BERT Devlin et al.
-(2018) and T5 (Raffel et al., 2020), the quality of text embedding can be substantially improved by pre-training with large-scale unlabeled data Izacard et al.
-(2021); Wang et al.
-(2022b).
+While one can resort to generic pre-trained models like BERT (Devlin et al., 2018) and T5 (Raffel et al., 2020), the quality of text embedding can be substantially improved by pre-training with large-scale unlabeled data (Izacard et al., 2021; Wang et al., 2022b).
 Further, instead of relying on a single algorithm, it takes a compound recipe to train general-purpose text embedding.
-Particularly, it needs embedding-oriented pre-training to prepare the text encoder Gao and Callan (2021), contrastive learning with sophisticated negative sampling to improve the embedding’s discriminability Qu et al.
-(2020), and instruction-based fine-tuning Su et al.
-(2022); Asai et al.
-(2022) to integrate different representation capabilities of text embedding.
+Particularly, it needs embedding-oriented pre-training to prepare the text encoder (Gao and Callan, 2021), contrastive learning with sophisticated negative sampling to improve the embedding’s discriminability (Qu et al., 2020), and instruction-based fine-tuning (Su et al., 2022; Asai et al., 2022) to integrate different representation capabilities of text embedding.
 $\bullet$ Benchmark.
 Another pre-requisite condition is the establishment of proper benchmarks, where all needed capabilities of text embeddings can be comprehensively evaluated.
-BEIR Thakur et al.
-(2021) provides a collection of 18 to evaluate the embedding’s general performances on different retrieval tasks, e.g., question answering and fact-checking.
-Later, MTEB Muennighoff et al.
-(2022a) proposes a more holistic evaluation of embeddings and extends BEIR.
+BEIR (Thakur et al., 2021) provides a collection of 18 to evaluate the embedding’s general performances on different retrieval tasks, e.g., question answering and fact-checking.
+Later, MTEB (Muennighoff et al., 2022a) proposes a more holistic evaluation of embeddings and extends BEIR.
 It integrates 56 datasets, where all important capabilities of text embeddings, like retrieval, ranking, clustering, etc., can be jointly evaluated.
 Altogether, the development of general-purpose text embedding needs to be made on top of a mixture of driving forces, from data, and encoder models, to training methods and benchmarking.
-In recent years, continual progress has been achieved in this field, such as work from Contriever Izacard et al.
-(2021), E5 Wang et al.
-(2022b), and OpenAI Text Embedding Neelakantan et al.
-(2022).
-However, most of these works are oriented to the English world.
-In contrast, there is a severe shortage of competitive models for general Chinese embedding due to a series of limitations: there are neither well-prepared training resources nor suitable benchmarks to evaluate the generality.
+In recent years, continual progresses have been achieved in this field, such as Contriever (Izacard et al., 2021), E5 (Wang et al., 2022b), GTR (Ni et al., 2021b), and OpenAI Text Embedding (Neelakantan et al., 2022).
+Nevertheless, most of these models are dedicated to the English-centric scenarios.
+In contrast, there is a shortage of competitive models for general Chinese embedding.
+What is worse, the development of general Chinese embedding is severely constrained in many aspects: there are neither well-prepared training resources nor suitable benchmarks to evaluate the generality.
+[^1] This situation has been substantially improved since our work.
+New methods are continually developed on top of the benchmark, data, and models from C-Pack.
 To address the above challenges, we present a package of resources called C-Pack, which contributes to the development of general Chinese embedding from the following perspectives.
@@
 [^2] https://huggingface.co/spaces/mteb/leaderboard C-MTEB collects 35 public-available datasets belonging to 6 types of tasks.
+We set up the unified testing protocols so that different embeddings can be evaluated on fair ground.
+We also develop the evaluation pipeline which significantly makes ease for the evaluation process.
 Thanks to the scale and diversity of C-MTEB, all major capabilities of Chinese embeddings can be reliably measured, making it the most suitable benchmark to evaluate the generality of Chinese text embedding.
 $\bullet$ C-MTP (Chinese Massive Text Pairs).
-We create a massive training dataset of 100M text pairs, which integrates both labeled data and unlabeled data curated from Wudao Yuan et al.
-(2021), one of the largest corpora for pre-training Chinese language models.
-C-MTP is not only large and diverse but also cleaned to ensure the data quality.
-$\bullet$ C-TEM (Chinese Text Embedding Models).
+We create a massive training dataset of 100M text pairs.
+The majority of our dataset is curated from the massive web corpora, such as Baike (Wikipedia-style webs in Chinese), Zhihu (a major Chinese social media), major News Websites in Chinese.
+We extract the semantically related text pairs leveraging the rich-structured information within the data, such as title-to-document, subtitle-to-passage, question-to-answer, question-to-similar-question, etc.
+The extracted data is further cleaned for the massive weakly supervised training of the text embeddings.
+We also integrate diverse labeled datasets, which presents high-quality supervision signals for the final refinement of text embeddings.
+Besides, considering that there is no public available dataset for general English text embeddings either, we curate another massive dataset for English with the same method, which consists of 200M text pairs.
+$\bullet$ BGE (BAAI General Embeddings).
 We provide a family of well-trained models for Chinese general text embeddings.
 There are three optional model sizes: small (24M), base (102M), and large (326M), which present users with the flexibility to trade off efficiency and effectiveness.
-Our models make a big leap forward in generality: C-TEM outperforms all previously Chinese text embedding models on all aspects of C-MTEB by large margins.
-Besides being directly applicable, C-TEM can also be fine-tuned with additional data for better task-specific performances.
+Our models make a big leap forward in generality: BGE outperforms all previously Chinese text embedding models on all aspects of C-MTEB by large margins.
+Besides being directly applicable, BGE can also be fine-tuned with additional data for better downstream performances.
+The releasing of these powerful models substantially contributes to critical applications, such as search, question answering, and retrieval-augmented generation.
 $\bullet$ Training Recipe.
@@
 The release of the training recipe will help the community to reproduce the state-of-the-art methods and make continuous progress on top of them.
-In summary, C-Pack provides a go-to option for people’s application of general-purpose Chinese text embedding.
-It substantially advances the training and evaluation, laying a solid foundation for the future development of this field.
-Figure 2: Overview of C-Pack.
+Our project enjoys a widespread popularity in technique communities, like HuggingFace and Github.
+Remarkably, according to the up-to-date statics (2024-04), BGE model series have received more than 20 million downloads from HuggingFace since its release on 2023-08, making it one of the most popular embedding models in the world.
+They have been integrated by the major RAG and text-embedding frameworks in the world, such as Langchain [^3] https://python.langchain.com/docs/integrations/text_embedding/bge_huggingface, LLamaIndex [^4] https://docs.llamaindex.ai/en/stable/examples/embeddings/huggingface.html, and Huggingface [^5] https://huggingface.co/docs/text-embeddings-inference/index.
+The code-base also receives nearly 5,000 stars on GitHub.
+So far, there have been over 100 submissions on C-MTEB, and it is widely recognized as the most popular and authoritative benchmark for Chinese text embeddings.
+It’s worth noting that our project is still fast growing, with new resources continually created and released to the public.
+In summary, C-Pack provides a go-to option for the development, evaluation, and application of general-purpose Chinese text embedding, which establishes a solid foundation for the advancement of this field.
+The remaining part of this paper is organized as follows.
+We discuss the related works about general text embeddings in Section 2.
+We present a detailed introduction about C-Pack resources in Section 3.
+We make comprehensive empirical analysis for the value of C-Pack in Section 4.
+Finally, we conclude our work in Section 5.
+Figure 2.
+Overview of C-Pack.
 C-MTEB is a benchmark for Chinese text embeddings.
 C-MTP is a large-scale Chinese embedding training dataset.
-C-TEM are state-of-the-art Chinese embedding models.
+BGE are state-of-the-art Chinese embedding models.
 The training recipe is shown at the bottom.
-2 C-Pack
-In this section, we first introduce the resources in C-Pack: the benchmark C-MTEB, the training data C-MTP, and the model class C-TEM.
+2. Related Work
+The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering (Karpukhin et al., 2020) but also due to its fundamental role in augmenting large language models (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Shi et al., 2023).
+Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios.
+In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever (Izacard et al., 2021), GTR (Ni et al., 2021b), sentence-T5 (Ni et al., 2021a), Sentence-Transformer (Reimers and Gurevych, 2019), E5 (Wang et al., 2022a), OpenAI text embedding (Neelakantan et al., 2022), etc.
+Although it remains an open problem, recent studies highlight the following important factors.
+$\bullet$ Firstly, the training data is desired to be large-scale and diversified, from which the embedding model can learn to recognize different kinds of semantic relationships (Izacard et al., 2021; Ni et al., 2021b; Wang et al., 2022b; Neelakantan et al., 2022).
+It usually calls for the comprehensive and elaborate data curation from web corpora, such as online encyclopedia, QA platforms, news websites, and social media communities.
+Despite similar efforts made by the previous works, few of the curated datasets are made public available before C-Pack.
+$\bullet$ Secondly, the embedding model must be scaled up in terms of both training and model size, as scaled-up text encoders are more generalizable across different application scenarios (Muennighoff, 2022; Ni et al., 2021b, a).
+Such an observation is in line with the conclusion for the importance of scaling LLMs (Hoffmann et al., 2022; Rae et al., 2021; Brown et al., 2020; Chowdhery et al., 2022; Srivastava et al., 2022; Gao et al., 2021a; Li et al., 2023a; Allal et al., 2023; Muennighoff et al., 2023b).
+Although large-scale training is affordable for many industrial organizations and companies, it remains a huge burden for the community users.
+The public release of BGE saves such a cost, benefiting both direct applications of text embeddings and further improvements for specific scenarios.
+$\bullet$ Thirdly, the training recipe must be optimized through pre-training (Liu and Shao, 2022; Wang et al., 2022a), negative sampling (Izacard et al., 2021; Wang et al., 2022a), and multi-task fine-tuning (Su et al., 2022; Asai et al., 2022; Sanh et al., 2021; Wei et al., 2021; Muennighoff et al., 2022b, 2023a; Chung et al., 2022).
+In C-Pack, these operations are integrated, optimized, and pipelined, which significantly facilitates people’s reproduction and continual fine-tuning of BGE.
+$\bullet$ Aside from the above factors, it is also critical to establish proper benchmarks to evaluate the generality of text embeddings.
+Unlike previous task-specific evaluations, like MSMARCO (Nguyen et al., 2016), SentEval (Conneau and Kiela, 2018), it is needed to substantially augment the benchmarks so as to evaluate the embedding’s performance for a wide variety of tasks.
+One representative work is made by BEIR (Thakur et al., 2021; Kamalloo et al., 2023), where the embeddings can be evaluated across different retrieval tasks.
+It is later extended by MTEB (Muennighoff et al., 2022a), where all major aspects of text embeddings can be comprehensively evaluated.
+However, no such works were done for the Chinese community before.
+By introducing C-MTEB, this limitation has been substantially conquered.
+Given the above analysis, it can be concluded that the general text embedding is highly resource-dependent, which calls for a wide range of elements, such as datasets, models, and benchmarks.
+Thus, the creation and public release of the corresponding resources in C-Pack is crucially important.
+3. C-Pack
+In this section, we first introduce the resources in C-Pack: the benchmark C-MTEB, the training data C-MTP, and the model class BGE.
 Then, we discuss the training recipe, which enables us to train the state-of-the-art models for general Chinese embedding based on the offered resources.
-2.1 Benchmark: C-MTEB
-[…]
-In particular, we collect a total of 35 public datasets, all of which can be used to evaluate Chinese text embeddings.
+3.1. Benchmark: C-MTEB
+C-MTEB is established for the comprehensive evaluation of the generality of Chinese embeddings (Figure 2).
+In the past few years, the community has put forward many datasets for text representation and language understanding tasks in Chinese, such as CMNLI (Xu et al., 2020), DuReader (He et al., 2017), T2Ranking (Xie et al., 2023).
+However, these datasets are independently curated, lacking a fair and shared ground to comprehensively evaluate the general capability of text embeddings.
+Therefore, we create C-MTEB, where the following important efforts are made: 1) the comprehensive collection of datasets which can be either directly utilized or repurposed for the evaluation of text embeddings, 2) the categorization of the datasets into different capability attributes of text embeddings, e.g., retrieval, similarity analysis, classification, etc., 3) the standardization of the evaluation protocols, 4) the establishment of evaluation pipelines.
+In particular, we collect a total of 35 public datasets related to the evaluation of Chinese text embeddings (Briefed as Figure 1.
+Detailed specifications are presented in the Github Repository [^6] https://github.com/FlagOpen/FlagEmbedding/tree/master/C_MTEB).
 The collected datasets are categorized based on the embedding’s capability they may evaluate.
@@
 The datasets of the same category are collected from different domains and complementary to each other, therefore ensuring the corresponding capability to be fully evaluated.
-The nature of each task and its evaluation metric are briefly introduced as follows.
+The nature of each evaluation task and the evaluation metric are briefly introduced as follows.
 $\bullet$ Retrieval.
@@
 The retrieval quality can be measured by ranking and recall metrics at different cut-offs.
-In this work, we use the setting from BEIR Thakur et al.
-(2021), using NDCG@10 as the main metric.
+In this work, we use the setting from BEIR (Thakur et al., 2021), using NDCG@10 as the main metric.
 $\bullet$ Re-ranking.
-The re-ranking task is presented with test queries and their lists of candidate documents (one positive plus N negative documents).
-For each query, it re-ranks the candidate documents based on the embedding similarity.
+The re-ranking task is presented with test queries and their candidate documents (1 positive plus N negative documents).
+For each query, it re-ranks the documents based on the embedding similarity.
 The MAP score is used as the main metric.
@@
 The STS (Agirre et al., 2012, 2013, 2014, 2015, 2016) task is to measure the correlation of two sentences based on their embedding similarity.
-Following the original setting in Sentence-BERT Reimers and Gurevych (2019), the Spearman’s correlation is computed with the given label, whose result is used as the main metric.
-$\bullet$ Classification.
-The classification task re-uses the logistic regression classifier from MTEB Muennighoff et al.
-(2022a), where the provided label is predicted based on the input embedding.
-The average precision is used as the main metric.
+Following the original setting in Sentence-BERT (Reimers and Gurevych, 2019), the Spearman’s correlation is computed with the given label, whose result is used as the main metric.
+$\bullet$ Classification. The classification task re-uses the logistic regression classifier from MTEB (Muennighoff et al., 2022a), where the average precision is used as the main metric.
 $\bullet$ Pair-classification.
@@
 The clustering task is to group sentences into meaningful clusters.
-Following the original setting in MTEB Muennighoff et al.
-(2022a), it uses the mini-batch k-means method for the evaluation, with batch size equal to 32 and k equal to the number of labels within the mini-batch.
+Following the original setting in MTEB (Muennighoff et al., 2022a), it uses the mini-batch k-means method for the evaluation, with batch size equal to 32 and k equal to the number of labels within the mini-batch.
 The V-measure score is used as the main metric.
+Figure 3. The evaluation pipeline of C-MTEB.
 Finally, the embedding’s capability on each task is measured by the average performance of all datasets for that task.
 The embedding’s overall generality is measured by the average performance of all datasets in C-MTEB.
-dataset
-C-MTP (unlabeled)
-C-MTP      (labeled)
-source
-Wudao, CSL, XLSUM-Zh, Amazon-Review-Zh, CMRC, etc.
-[…]
-size
-[…]
-[…]
-Table 1: Composition of C-MTP
-2.2 Training Data: C-MTP
+We set up the standardized and compact pipeline for all tasks where different embedding models can be evaluated on a fair basis (Figure 3).
+FlagDRESModel is the wapper for the custmoized embedding model, which implements the encoding method for query and document.
+ChineseTaskList is the task list of C-MTEB.
+The evaluation result is written to the output folder, which can be directly submitted to the C-MTEB leaderboard [^7] https://huggingface.co/spaces/mteb/leaderboard.
+3.2. Training Data: C-MTP
 We curate the largest dataset C-MTP for the training of general Chinese embedding.
@@
 To ensure the generality of the text embedding, the paired texts need to be both large-scale and diversified.
-Therefore, C-MTP is collected from two sources: the curation of massive unlabeled data, a.k.a.
-C-MTP (unlabeled); and the comprehensive collection of labeled data, a.k.a.
-C-MTP (labeled).
+Therefore, C-MTP is collected from two sources.
+The majority of the data is based on the curation of massive unlabeled data, a.k.a.
+C-MTP (unlabeled), which presents 100 millions of paired texts.
+Meanwhile, a small portion is from the comprehensive integration of high-quality labeled data, a.k.a.
+C-MTP (labeled), which leads to about 1 million paired texts.
 The data collection process is briefly introduced as follows.
+Table 1. Composition of C-MTP
+dataset C-MTP (unlabeled) C-MTP      (labeled)
+source Wudao, Zhihu, Baike, CSL, XLSUM-Zh, Amazon-Review-Zh, CMRC, etc. T2-Ranking, mMARCO-Zh, DuReader, NLI-Zh, etc.
+[…]
+Figure 4. Creation of C-MTP.
 $\bullet$ C-MTP (unlabeled).
@@
 Our primary source of data comes from open web corpora.
-The most representative one is the Wudao corpus Yuan et al.
-(2021), which is the largest well-formatted dataset for pre-training Chinese language models.
-For each of its articles, we extract (title, passage) to form a text pair.
-Following the same recipe, we also collect such text pairs from other similar web content like Zhihu, Baike, news websites, etc.
-Aside from the open web content, we also explore other public Chinese datasets to extract text pairs, such as CSL (scientific literature), Amazon-Review-Zh (reviews), Wiki Atomic Edits (paraphrases), CMRC (machine reading comprehension), XLSUM-Zh (summarization), etc.
-The paired structures are obvious in these datasets, which are directly extracted for the augmentation of C-MTP (unlabeled).
+The most representative one is the Wudao corpus (Yuan et al., 2021), which is the largest dataset of well-formatted articles for pre-training Chinese language models.
+For each its article, we extract structures, like (title, body), (sub-title, passage), to form a text pair.
+In addition, we collect data from other web content like Zhihu, Baike, news websites, which complement other forms of text pairs, especially (question, answer), (paraphrase titles), (paraphrase answers), etc.
+Aside from the open web content, we also explore other public Chinese data for more diverse text pairs, such as CSL (scientific literature), Amazon-Review-Zh (topic and its reviews), Wiki Atomic Edits (paraphrases), CMRC (machine reading comprehension), XLSUM-Zh (summarization), etc.
 The text pairs curated from the web and other public sources are not guaranteed to be closely related.
 Therefore, data quality can be a major concern.
-In our work, we use a simple strategy to filter the data before adding it to C-MTP (unlabeled).
-Particularly, we use a third-party model: Text2Vec-Chinese [^2] https://huggingface.co/GanymedeNil to score the strength of relation for each text pair.
+In our work, we make use of a compound data cleaning strategy to refine the raw data.
+Firstly, the whole data undergoes general filtering, which removes non-textual, duplicated, and malicous content.
+Secondly, the data is further processed by semantic filtering so as to ensure the text pairs are semantically related.
+In our work, we make use a third-party model: Text2Vec-Chinese [^8] https://huggingface.co/GanymedeNil to score the strength of relation for each text pair.
 We empirically choose a threshold of 0.43, and drop the samples whose scores are below the threshold.
@@
 Despite the simplicity, we find that it effectively removes the irrelevant text pairs when manually reviewing samples and leads to strong empirical performances for the models trained on C-MTP (unlabeled).
-[…]
+$\bullet$ C-MTP (labeled).
+The labeled data is collected to further enhance to the training data.
+In our work, we integrate and re-purpose a diverse group of datasets, which covers different capabilities of the text embedding, like retrieval, ranking, similarity comparison, etc.
+Particularly, the following labeled datasets are included, T2-Ranking (Xie et al., 2023), DuReader (He et al., 2017; Qiu et al., 2022), mMARCO (Bonifacio et al., 2021), CMedQA-v2(Zhang et al., 2018), multi-cpr(Long et al., 2022), NLI-Zh [^9] https://huggingface.co/datasets/shibing624/nli_zh, cmnli(Xu et al., 2020) and ocnli(Xu et al., 2020).
+There are 838,465 paired texts in total, which contains diverse question-answering and paraphrasing patterns.
+Although it is much smaller than C-MTP (unlabeled), most of the data is curated from human annotation, thus ensuring a high credibility of relevance.
 Given the differences in scale and quality, C-MTP (unlabeled) and C-MTP (labeled) are applied to different training stages, which jointly result in a strong performance for the embedding model.
 Detailed analysis will be made in our training recipe.
-2.3 Model Class: C-TEM
-We provide a comprehensive class of well-trained embedding models for the community.
-Our models take a BERT-like architecture, where the last layer’s hidden state of the special token [CLS] is trained to work as the embedding.
-There are three different scales for the models: large (with 326M parameters), base (with 102M parameters), and small (with 24M parameters).
+3.3. Model Class: BGE
+Even with the full package of training data, it is still challenging to learn general text embeddings due to the expensive training process.
+In our work, we provide a comprehensive class of well-trained embedding models for the community.
+Our models are based on the BERT-like architecture (Devlin et al., 2018), which go through three-stage of training (to be discussed in the next section).
+There are three available scales: large (with 326M parameters), base (with 102M parameters), and small (with 24M parameters).
 The large-scale model achieves the highest general representation performances, leading the current public-available models by a considerable margin.
-The small-scale model is also empirically competitive compared with the public-available models and other model options in C-TEM; besides, it is way faster and lighter, making it suitable to handle massive knowledge bases and high-throughput applications.
+The small-scale model is also empirically competitive compared with the public-available models and other model options in BGE; besides, it is way faster and lighter, making it suitable to handle massive knowledge bases and high-throughput applications.
 Thanks to the comprehensive coverage of different model sizes, people are presented with the flexibility to trade off running efficiency and representation quality based on their own needs.
-As introduced, the models within C-TEM have been well-trained and achieve a strong generality for a wide variety of tasks.
-Meanwhile, they can also be further fine-tuned if 1) the embeddings are applied for a specific scenario, 2) the training data is presented for the application scenario.
-It is empirically verified that the fine-tuned model may bring forth a much better performance for its application, compared with its original model in C-TEM, and the fine-tuned models from other general pre-trained encoders, like BERT.
-In other words, C-TEM not only presents people with direct usage embeddings but also works as a foundation where people may develop more powerful embeddings.
-2.4 Training Recipe
-The training recipe of C-TEM is completely released to the public along with C-Pack (Figure 2).
-Our training recipe has three main components: 1) pre-training with plain texts, 2) contrastive learning with C-MTP (unlabeled), and 3) multi-task learning with C-MTP (labeled), whose specifications are made as follows.
+As introduced, the models within BGE have been well-trained and achieve a strong generality for a wide variety of tasks.
+Meanwhile, they also establish a strong foundation for further fine-tuning.
+The fine-tuning recipe is well formulated in our training recipe, where all people’s need is the preparation of fine-tuning data.
+It is empirically verified that the fine-tuned model may bring forth a much better performance for its application, compared with its original model in BGE, and the fine-tuned models from other general pre-trained encoders, like BERT.
+In other words, BGE not only presents people with direct usage embeddings but also works as a foundation where people may develop more powerful embeddings.
+Table 2. Performance of various models on C-MTEB.
+model Dim Retrieval STS Pair CLF CLF Re-rank Cluster Average
+Text2Vec (base) 768 38.79 43.41 67.41 62.19 49.45 37.66 48.59
+Text2Vec (large) 1024 41.94 44.97 70.86 60.66 49.16 30.02 48.56
+Luotuo (large) 1024 44.40 42.79 66.62 61.0 49.25 44.39 50.12
+M3E (base) 768 56.91 50.47 63.99 67.52 59.34 47.68 57.79
+M3E (large) 1024 54.75 50.42 64.30 68.20 59.66 48.88 57.66
+Multi. E5 (base) 768 61.63 46.49 67.07 65.35 54.35 40.68 56.21
+Multi. E5 (large) 1024 63.66 48.44 69.89 67.34 56.00 48.23 58.84
+OpenAI-Ada-002 1536 52.00 43.35 69.56 64.31 54.28 45.68 53.02
+BGE (small) 512 63.07 49.45 70.35 63.64 61.48 45.09 58.28
+BGE (base) 768 69.53 54.12 77.50 67.07 64.91 47.63 62.80
+BGE (large) 1024 71.53 54.98 78.94 68.32 65.11 48.39 63.96
+3.4. Training Recipe
+The training recipe of BGE is completely released to the public along with C-Pack (Figure 2).
+Our training recipe has three main components: 1) pre-training with plain texts, 2) contrastive learning with C-MTP (unlabeled), and 3) multi-task learning with C-MTP (labeled).
+As introduced, the public release of training recipe will not only help with the reproduction, but also benefit the improvement of BGE with continual training and fine-tuning.
 $\bullet$ Pre-Training.
 Our model is pre-trained on massive plain texts through a tailored algorithm in order to better support the embedding task.
-Particularly, we make use of the Wudao corpora Yuan et al.
-(2021), which is a huge and high-quality dataset for Chinese language model pre-training.
-We leverage the MAE-style approach presented in RetroMAE Liu and Shao (2022); Xiao et al.
-(2023), which is simple but highly effective.
+Particularly, we make use of the Wudao corpora (Yuan et al., 2021), which is a huge and high-quality dataset for Chinese language model pre-training.
+We leverage the MAE-style approach presented in RetroMAE (Liu and Shao, 2022; Xiao et al., 2023), which is simple but highly effective.
 The polluted text is encoded into its embedding, from which the clean text is recovered on top of a light-weight decoder:
 $\min.\sum_{x\in X}-\log\mathrm{Dec}(x|\mathbf{e}_{\tilde{X}}),~{}\mathbf{e}_{\tilde{X}}\leftarrow\mathrm{Enc}(\tilde{X}).$
-($\mathrm{Enc}$, $\mathrm{Dec}$ are the encoder and decoder, $X$, $\tilde{X}$ indicate the clean and polluted text.)
+($\mathrm{Enc}$, $\mathrm{Dec}$ indicate the encoding and decoding operations, $X$, $\tilde{X}$ indicate the clean and polluted text.)
 $\bullet$ General purpose fine-tuning.
@@
 One critical factor of contrastive learning is the negative samples.
-Instead of mining hard negative samples on purpose, we purely rely on in-batch negative samples Karpukhin et al.
-(2020) and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
-model
-Dim
-Retrieval
-STS
-Pair CLF
-CLF
-Re-rank
-Cluster
-Average
-Text2Vec (base)
-768
-38.79
-43.41
-67.41
-62.19
-49.45
-37.66
-48.59
-Text2Vec (large)
-1024
-41.94
-44.97
-70.86
-60.66
-49.16
-30.02
-48.56
-Luotuo (large)
-1024
-44.40
-42.79
-66.62
-61.0
-49.25
-44.39
-50.12
-M3E (base)
-768
-56.91
-50.47
-63.99
-67.52
-59.34
-47.68
-57.79
-M3E (large)
-1024
-54.75
-50.42
-64.30
-68.20
-59.66
-48.88
-57.66
-Multi. E5 (base)
-768
-61.63
-46.49
-67.07
-65.35
-54.35
-40.68
-56.21
-Multi. E5 (large)
-1024
-63.66
-48.44
-69.89
-67.34
-56.00
-48.23
-58.84
-OpenAI-Ada-002
-1536
-52.00
-43.35
-69.56
-64.31
-54.28
-45.68
-53.02
-BGE (small)
-512
-63.07
-49.45
-70.35
-63.64
-61.48
-45.09
-58.28
-BGE (base)
-768
-69.53
-54.12
-77.50
-67.07
-64.91
-47.63
-62.80
-BGE (large)
-1024
-71.53 54.98 78.94 68.32 65.11
-48.39
-63.96
-Table 2: Performance of various models on C-MTEB.
+Instead of mining hard negative samples on purpose, we purely rely on in-batch negative samples (Karpukhin et al., 2020) and resort to a big batch size (as large as 19,200) to improve the discriminativeness of the embedding.
 $\bullet$ Task-specific fine-tuning.
@@
 In this place, we apply two strategies to mitigate this problem.
-On one hand, we leverage instruction-based fine-tuning Su et al.
-(2022); Asai et al.
-(2022), where the input is differentiated to help the model accommodate different tasks.
+On one hand, we leverage instruction-based fine-tuning (Su et al., 2022; Asai et al., 2022), where the input is differentiated to help the model accommodate different tasks.
 For each text pair ($p$, $q$), a task specific instruction $I_{t}$ is attached to the query side: $q^{\prime}\leftarrow q+I_{t}$.
@@
 On the other hand, the negative sampling is updated: in addition to the in-batch negative samples, one hard negative sample $q^{\prime}$ is mined for each text pair ($p$, $q$).
-The hard negative sample is mined from the task’s original corpus, following the ANN-style sampling strategy in Xiong et al.
-(2020).
-3 Experiments
-We conduct experimental studies for the following purposes.
+The hard negative sample is mined from the task’s original corpus, following the ANN-style sampling strategy in (Xiong et al., 2020).
+4. Experiments
+In this section, we conduct experimental studies for the exploration of following problems.
 P1.
@@
 P2.
-The empirical verification of the text embeddings by C-TEM.
+The empirical verification of the text embeddings by BGE.
 P3.
@@
 The exploration of the impacts introduced by the training recipe.
-We consider the following popular Chinese text embedding models as the baselines for our experiments: Text2Vec-Chinese [^4] https://huggingface.co/shibing624 base and large; Luotuo [^5] https://huggingface.co/silk-road/luotuo-bert-medium; M3E [^6] https://huggingface.co/moka-ai base and large; multilingual E5 Wang et al.
-(2022b) and OpenAI text embedding ada 002777https://platform.openai.com/docs/guides/embeddings.
-The main metric presented in Section 2.1 is reported for each task in C-MTEB.
-3.1 General Evaluation
-We extensively evaluate C-TEM against popular Chinese text embeddings on C-MTEB as shown in Table 2.888Our C-TEM models are named BGE in the tables. We make the following observations.
+We consider the following popular Chinese text embedding models as the baselines for our experiments: Text2Vec-Chinese [^10] https://huggingface.co/shibing624 base and large; Luotuo [^11] https://huggingface.co/silk-road/luotuo-bert-medium; M3E [^12] https://huggingface.co/moka-ai base and large; multilingual E5 (Wang et al., 2022b) and OpenAI text embedding ada 002131313https://platform.openai.com/docs/guides/embeddings.
+The main metric presented in Section 3.1 is reported for each evaluation task in C-MTEB.
+Table 3. Ablation of the training data, C-MTP, and the training recipe.
+model Dim Retrieval STS Pair CLF CLF Re-rank Cluster Average
+M3E (large) 1024 54.75 50.42 64.30 68.20 59.66 48.88 57.66
+OpenAI-Ada-002 1536 52.00 43.35 69.56 64.31 54.28 45.68 53.02
+BGE-pretrain 1024 63.90 47.71 61.67 68.59 60.12 47.73 59.00
+BGE w.o. pre-train 1024 62.56 48.06 61.66 67.89 61.25 46.82 58.62
+BGE w.o. Instruct 1024 70.55 53.00 76.77 68.58 64.91 50.01 63.40
+BGE-finetune 1024 71.53 54.98 78.94 68.32 65.11 48.39 63.96
+4.1. General Evaluation
+We extensively evaluate BGE against popular Chinese text embeddings on C-MTEB as shown in Table 2.141414Our BGE models are named BGE in the tables., where we can make the following observations.
 First, our models outperform existing Chinese text embeddings by large margins.
@@
 Although the advantages for classification and clustering tasks are not as obvious, our performances are still on par or slightly better than the other most competitive models.
-The above observations verify the strong generality of C-TEM.
+The above observations verify the strong generality of BGE.
 Our models can be directly utilized to support different types of application scenarios.
@@
 As a result, it provides people with the flexibility to trade-off embedding quality and running efficiency: people may resort to our large-scale embedding model to deal with high-precision usages, or switch to the small-scale one for high-throughput scenarios.
-model
-Dim
-Retrieval
-STS
-Pair CLF
-CLF
-Re-rank
-Cluster
-Average
-M3E (large)
-1024
-54.75
-50.42
-64.30
-68.20
-59.66
-48.88
-57.66
-OpenAI-Ada-002
-1536
-52.00
-43.35
-69.56
-64.31
-54.28
-45.68
-53.02
-w.o. Instruct
-1024
-70.55
-53.00
-76.77
-68.58
-64.91
-50.01
-63.40
-BGE-$i$
-1024
-63.90
-47.71
-61.67
-68.59
-60.12
-47.73
-59.00
-BGE-$i$ w.o. pre-train
-1024
-62.56
-48.06
-61.66
-67.89
-61.25
-46.82
-58.62
-BGE-$f$
-1024
-71.53 54.98 78.94
-68.32
-65.11
-48.39
-63.96
-Table 3: Ablation of the training data, C-MTP, and the training recipe.
-3.2 Detailed Analysis
-We investigate the detailed impact of C-MTP and our training recipe. The corresponding experiment results are presented in Table 3 and Table 4.
+Using the same recipe as the Chinese models, we also train a set of English text embedding models presented in Table 5.
+Besides, our English data was created and released together with C-MTP.
+It was the first time that such comprehensive training data was made publicly available.
+At the time of public release, our English BGE models achieved the state-of-the-art performance on the English MTEB benchmark (Muennighoff et al., 2022a) across its 56 datasets.
+Although there were many strong competitors in the English community, such as E5 (Wang et al., 2022b), SGPT (Muennighoff, 2022), GTE (Li et al., 2023b), GTR (Ni et al., 2021b), and OpenAI Ada-002 (Neelakantan et al., 2022), we were able to notably advance the prior SOTA by an absolute 1.1 points in total average, which further verify the effectiveness of our data curation and training method.
+4.2. Detailed Analysis
+We investigate the detailed impact of C-MTP and our training recipe. The corresponding experiment results are presented in Table 3 and Table 4, respectively.
 First of all, we analyze the impact of our training data, C-MTP.
 As mentioned, C-MTP consists of two parts.
-1) C-MTP (unlabeled), which is used for general-purpose fine-tuning; the model produced from this stage is called the intermediate checkpoint, denoted as BGE-$i$.
-2) C-MTP (labeled), where the task-specific fine-tuning is further conducted on top of BGE-$i$; the model produced from this stage is called the final checkpoint, noted as BGE-$f$.
+1) C-MTP (unlabeled), which is used for general-purpose fine-tuning; the model produced from this stage is called the intermediate checkpoint, denoted as BGE-pretrain.
+2) C-MTP (labeled), where the task-specific fine-tuning is further conducted on top of BGE-pretrain; the model produced from this stage is called the final checkpoint, noted as BGE-finetune.
 Based on our observations from the experimental result, both C-MTP (unlabeled) and C-MTP (labeled) substantially contribute to the embedding’s quality.
 Regarding C-MTP (unlabeled), despite mostly being curated from unlabeled corpora, this dataset alone brings forth strong empirical performance for the embedding models trained on it.
-Compared with other baselines like Text2Vec, M3E, and OpenAI text embedding, BGE-$i$ already achieves a higher average performance.
+Compared with other baselines like Text2Vec, M3E, and OpenAI text embedding, BGE-pretrain already achieves a higher average performance.
 A further look into the performances reveals more details.
-On one hand, C-MTP (unlabeled) makes a major impact on the embedding’s retrieval quality, where BGE-$i$ notably outperforms the baselines in this attribute.
-On the other hand, the general capability of embedding is primarily established with C-MTP (unlabeled), as BGE-$i$’s performance is close to the baselines on the rest of the aspects, like STS and Clustering.
+On one hand, C-MTP (unlabeled) makes a major impact on the embedding’s retrieval quality, where BGE-pretrain notably outperforms the baselines in this attribute.
+On the other hand, the general capability of embedding is primarily established with C-MTP (unlabeled), as BGE-pretrain’s performance is close to the baselines on the rest of the aspects, like STS and Clustering.
 This puts our embedding models in a very favorable position for further improvements.
 As for C-MTP (labeled), the dataset is much smaller but of better quality.
-With another round of fine-tuning on C-MTP (labeled), the empirical advantage is significantly expanded for the final checkpoint BGE-$f$, where it gives rise to a jump in average performance from 59.0 (BGE-$i$) to 63.96 (BGE-$f$).
+With another round of fine-tuning on C-MTP (labeled), the empirical advantage is significantly expanded for the final checkpoint BGE-finetune, where it gives rise to a jump in average performance from 59.0 (BGE-pretrain) to 63.96 (BGE-finetune).
 Knowing that the text pairs in C-MTP (labeled) are mainly gathered from retrieval and NLI tasks, the most notable improvements are achieved on closely related tasks, namely retrieval, re-ranking, STS, and pair classification.
@@
 This indicates that a mixture of high-quality and diversified labeled data is able to bring forth substantial and comprehensive improvements for a pre-trained embedding model.
-We further explore the impact of our training recipe, particularly contrastive learning, task-specific fine-tuning, and pre-training.
+Table 4. Impact of batch size.
+Task Batch Size 256 2,048 19,200
+Retrieval 57.25 60.96 63.90
+STS 46.16 46.60 47.71
+Pair CLF 62.02 61.91 61.67
+CLF 65.71 67.42 68.59
+Re-rank 58.59 59.98 60.12
+Cluster 49.52 49.04 47.73
+Average 56.43 57.92 59.00
+Table 5. Performance of English Models on MTEB.
+Model Name Dim. Average Retrieval Cluster Pair CLF Re-rank STS Summarize CLF
+GTE (large) 1024 63.13 52.22 46.84 85.00 59.13 83.35 31.66 73.33
+GTE (base) 768 62.39 51.14 46.2 84.57 58.61 82.3 31.17 73.01
+E5 (large) 1024 62.25 50.56 44.49 86.03 56.61 82.05 30.19 75.24
+Instructor-XL 768 61.79 49.26 44.74 86.62 57.29 83.06 32.32 61.79
+E5 (base) 768 61.5 50.29 43.80 85.73 55.91 81.05 30.28 73.84
+GTE (small) 384 61.36 49.46 44.89 83.54 57.7 82.07 30.42 72.31
+OpenAI Ada 002 1536 60.99 49.25 45.9 84.89 56.32 80.97 30.8 70.93
+E5 (small) 384 59.93 49.04 39.92 84.67 54.32 80.39 31.16 72.94
+ST5 (XXL) 768 59.51 42.24 43.72 85.06 56.42 82.63 30.08 73.42
+MPNet (base) 768 57.78 43.81 43.69 83.04 59.36 80.28 27.49 65.07
+SGPT Bloom (7.1B) 4096 57.59 48.22 38.93 81.9 55.65 77.74 33.60 66.19
+BGE (small) 384 62.17 51.68 43.82 84.92 58.36 81.59 30.12 74.14
+BGE (base) 768 63.55 53.25 45.77 86.55 58.86 82.4 31.07 75.53
+BGE (large) 1024 64.23 54.29 46.08 87.12 60.03 83.11 31.61 75.97
+We make further exploration about our training recipe, particularly the impact from contrastive learning, task-specific fine-tuning, and pre-training.
 One notable feature of our training recipe is that we adopt a large batch size for contrastive learning.
-According to previous studies, the learning of the embedding model may benefit from the increasing of negative samples Izacard et al.
-(2021); Qu et al.
-(2020); Muennighoff (2022).
+According to previous studies, the learning of the embedding model may benefit from the increasing of negative samples (Izacard et al., 2021; Qu et al., 2020; Muennighoff, 2022).
 Given our dependency on in-batch negative samples, the batch size needs to be expanded as much as possible.
-In our implementation, we use a compound strategy of gradient checkpointing and cross-device embedding sharing Gao et al.
-(2021b), which results in a maximum batch size of 19,200.
+In our implementation, we use a compound strategy of gradient checkpointing and cross-device embedding sharing (Gao et al., 2021b), which results in a maximum batch size of 19,200.
 By making a parallel comparison between bz: 256, 2028, 19,200, we observe consistent improvement in embedding quality with the expansion of batch size (noted as bz).
@@
 C-MTP (labeled), where the model is fine-tuned on a small group of tasks.
-This indicates that using instructions may substantially contribute to task-specific fine-tuning.
-One more characteristic is that we use a specifically pre-trained text encoder to train C-TEM, rather than using common choices, like BERT and RoBERTa Liu et al.
-(2019).
-To explore its impact, we replace the pre-trained text encoder with the widely used Chinese-RoBERTa999huggingface.co/hfl/chinese-roberta-wwm-ext-large, noted as “BGE-$i$ w.o.
+This indicates that using instructions may substantially contribute to the quality of task-specific fine-tuning.
+One more characteristic is that we use a specifically pre-trained text encoder to train BGE, rather than using common choices, like BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019).
+To explore its impact, we replace the pre-trained text encoder with the widely used Chinese-RoBERTa151515huggingface.co/hfl/chinese-roberta-wwm-ext-large, noted as “BGE w.o.
 pre-train”.
-According to the comparison with BGE-$i$, the pre-trained text encoder notably improves the retrieval capability, while preserving similar performances on other aspects.
-Batch Size
-256
-2,048
-19,200
-Retrieval
-57.25
-60.96
-63.90
-STS
-46.16
-46.60
-47.71
-Pair CLF
-62.02
-61.91
-61.67
-CLF
-65.71
-67.42
-68.59
-Re-rank
-58.59
-59.98
-60.12
-Cluster
-49.52
-49.04
-47.73
-Average
-56.43
-57.92
-59.00
-Table 4: Impact of batch size.
-4 Related Work
-The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering Karpukhin et al.
-(2020) but also due to its fundamental role in augmenting large language models Lewis et al.
-(2020); Guu et al.
-(2020); Borgeaud et al.
-(2022); Izacard et al.
-(2022); Shi et al.
-(2023).
-Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios.
-In recent years, there has been a continual effort in this field, where a series of well-known works are proposed, like Contriever Izacard et al.
-(2021), GTR Ni et al.
-(2021b), sentence-T5 Ni et al.
-(2021a), Sentence-Transformer Reimers and Gurevych (2019), E5 Wang et al.
-(2022a), OpenAI text embedding Neelakantan et al.
-(2022), etc.
-Although it remains an open problem, recent studies highlight the following important factors.
-Firstly, the training data is desired to be large-scale and diversified, from which the embedding model can learn to recognize different kinds of semantic relationships Izacard et al.
-(2021); Wang et al.
-(2022b); Neelakantan et al.
-(2022).
-Secondly, the embedding model must be scaled up, as large text encoders are more generalizable across different application scenarios Muennighoff (2022); Ni et al.
-(2021b, a) in line with observations for the importance of scaling LLMs (Hoffmann et al., 2022; Rae et al., 2021; Brown et al., 2020; Chowdhery et al., 2022; Srivastava et al., 2022; Gao et al., 2021a; Li et al., 2023a; Allal et al., 2023; Muennighoff et al., 2023b).
-Thirdly, the training recipe must be optimized through pre-training Liu and Shao (2022); Wang et al.
-(2022a), negative sampling Izacard et al.
-(2021); Wang et al.
-(2022a), and multi-task fine-tuning Su et al.
-(2022); Asai et al.
-(2022); Sanh et al.
-(2021); Wei et al.
-(2021); Muennighoff et al.
-(2022b, 2023a); Chung et al.
-(2022).
-Aside from the above, it is also critical to establish proper benchmarks to evaluate the generality of text embeddings.
-Unlike previous task-specific evaluations, like MSMARCO Nguyen et al.
-(2016), SentEval Conneau and Kiela (2018), it is needed to substantially augment the benchmarks so as to evaluate the embedding’s performance for a wide variety of tasks.
-One representative work is made by BEIR Thakur et al.
-(2021); Kamalloo et al.
-(2023), where the embeddings can be evaluated across different retrieval tasks.
-It is later extended by MTEB Muennighoff et al.
-(2022a), where all major aspects of text embeddings can be comprehensively evaluated.
-Given the above analysis, it can be concluded that the general text embedding is highly resource-dependent, which calls for a wide range of elements, such as datasets, models, and benchmarks.
-Thus, the creation and public release of the corresponding resources is crucially important.
-5 Conclusion
-We present C-Pack to advance progress towards general Chinese embedding.
-C-Pack consists of three core resources 1) The benchmark C-MTEB, covering 6 major tasks of embeddings and 35 datasets, making it the most comprehensive benchmark to evaluate the generality of Chinese embeddings.
-2) The training data C-MTP, curated from massive unlabeled corpora and high-quality labeled datasets.
-Its unprecedented scale, diversity, and quality contribute to the superior generality of our embedding models.
-3) The models C-TEM, which are empirically competitive.
-Their different sizes provide people with the flexibility to trade off efficiency and embedding quality.
-The entire training recipe is also provided along with these resources.
-The public release of C-Pack facilitates the usage of general Chinese embedding and also paves the way for its future advancement.
-Appendix A C-MTEB Datasets
-Table 5: Overview of datasets in C-MTEB.
-Name URL Description Categ. Test        Samples
-Classification
-AmazonReviewsClassification Muennighoff et al. (2022a); McAuley and Leskovec (2013)
-https://hf.co/datasets/mteb/amazon_reviews_multi
-Sentiment of Amazon reviews
-s2s
-5,000
-IFlyTek Xu et al. (2020a)
-https://hf.co/datasets/C-MTEB/IFlyTek-classification
-Long text classification for App descriptions
-s2s
-2,600
-JDReview (https://hf.co/datasets/kuroneko5943/jd21)
-https://hf.co/datasets/C-MTEB/JDReview-classification
-iPhone reviews
-s2s
-533
-MassiveIntentClassification Muennighoff et al. (2022a); FitzGerald et al. (2022)
-https://hf.co/datasets/mteb/amazon_massive_intent
-Amazon Alexa virtual assistant utterances annotated with the associated intent
-s2s
-16,500
-MassiveScenarioClassification Muennighoff et al. (2022a); FitzGerald et al. (2022)
-https://hf.co/datasets/mteb/amazon_massive_scenario
-Amazon Alexa virtual assistant utterances annotated with the associated scenario
-s2s
-16,500
-MultilingualSentiment McAuley and Leskovec (2013)
-https://hf.co/datasets/C-MTEB/MultilingualSentiment-classification
-Sentiment of Amazon reviews
-s2s
-3,000
-OnlineShopping (https://github.com/SophonPlus/ChineseNlpCorpus/blob/master/datasets/online_shopping_10_cats/intro.ipynb)
-https://hf.co/datasets/C-MTEB/OnlineShopping-classification
-Sentiment Analysis of User Reviews on Online Shopping Websites
-s2s
-1,000
-TNews Xu et al. (2020a)
-https://hf.co/datasets/C-MTEB/TNews-classification
-Short Text Classification for News
-s2s
-10,000
-Waimai (https://github.com/SophonPlus/ChineseNlpCorpus/blob/master/datasets/waimai_10k/intro.ipynb)
-https://hf.co/datasets/C-MTEB/waimai-classification
-Sentiment Analysis of user reviews on takeaway platforms
-s2s
-1,000
-Clustering
-CLSClusteringP2P Li et al. (2022)
-https://hf.co/datasets/C-MTEB/CLSClusteringP2P
-Clustering of titles + abstract from CLS dataset. Clustering of 13 sets, based on the main category.
-p2p
-10,000
-CLSClusteringS2S Li et al. (2022)
-https://hf.co/datasets/C-MTEB/C-MTEB/CLSClusteringS2S
-Clustering of titles from CLS dataset. Clustering of 13 sets, based on the main category.
-s2s
-10,000
-ThuNewsClusteringP2P Li et al. (2006); Li and Sun (2007)
-https://hf.co/datasets/C-MTEB/ThuNewsClusteringP2P
-Clustering of titles + abstract from the THUCNews dataset
-p2p
-10,000
-ThuNewsClusteringS2S Li et al. (2006); Li and Sun (2007)
-https://hf.co/datasets/C-MTEB/ThuNewsClusteringS2S
-Clustering of titles from the THUCNews dataset
-s2s
-10,000
-Pair Classification
-Cmnli Xu et al. (2020a, b); Conneau and Kiela (2018); Williams et al. (2017)
-https://hf.co/datasets/C-MTEB/CMNLI
-Chinese Multi-Genre NLI
-s2s
-139,000
-Ocnli Hu et al. (2020)
-https://hf.co/datasets/C-MTEB/OCNLI
-Original Chinese Natural Language Inference dataset
-s2s
-3,000
-Reranking
-T2Reranking Xie et al. (2023)
-https://hf.co/datasets/C-MTEB/T2Reranking
-T2Ranking: A large-scale Chinese Benchmark for Passage Ranking
-s2p
-24,382
-MMarcoRetrieval Bonifacio et al. (2021)
-https://hf.co/datasets/C-MTEB/Mmarco-reranking
-mMARCO is a multilingual version of the MS MARCO passage ranking dataset
-s2p
-7,437
-CMedQAv1 Zhang et al. (2017)
-https://hf.co/datasets/C-MTEB/CMedQAv1-reranking
-Chinese community medical question answering
-s2p
-2,000
-CMedQAv2 Zhang et al. (2018b)
-https://hf.co/datasets/C-MTEB/C-MTEB/CMedQAv2-reranking
-Chinese community medical question answering
-s2p
-4,000
-Retrieval
-T2Retrieval Xie et al. (2023)
-https://hf.co/datasets/C-MTEB/T2Retrieval
-T2Ranking: A large-scale Chinese Benchmark for Passage Ranking
-s2p
-24,832
-MMarcoRetrieval Bonifacio et al. (2021)
-https://hf.co/datasets/C-MTEB/MMarcoRetrieval
-mMARCO is a multilingual version of the MS MARCO passage ranking dataset
-s2p
-7,437
-DuRetrieval Qiu et al. (2022)
-https://hf.co/datasets/C-MTEB/DuRetrieval
-A Large-scale Chinese Benchmark for Passage Retrieval from Web Search Engine
-s2p
-4,000
-CovidRetrieval Qiu et al. (2022)
-https://hf.co/datasets/C-MTEB/CovidRetrieval
-COVID-19 news articles
-s2p
-949
-CmedqaRetrieval Qiu et al. (2022)
-https://hf.co/datasets/C-MTEB/CmedqaRetrieval
-Online medical consultation text
-s2p
-3,999
-EcomRetrieval Long et al. (2022)
-https://hf.co/datasets/C-MTEB/EcomRetrieval
-Passage retrieval dataset collected from Alibaba search engine systems in e-commerce domain
-s2p
-1,000
-MedicalRetrieval Long et al. (2022)
-https://hf.co/datasets/C-MTEB/MedicalRetrieval
-Passage retrieval dataset collected from Alibaba search engine systems in medical domain
-s2p
-1,000
-VideoRetrieval Long et al. (2022)
-https://hf.co/datasets/C-MTEB/VideoRetrieval
-Passage retrieval dataset collected from Alibaba search engine systems in video domain
-s2p
-1,000
-STS
-AFQMC Xu et al. (2020a)
-https://hf.co/datasets/C-MTEB/AFQMC
-Ant Financial Question Matching Corpus
-s2s
-3,861
-ATEC (https://github.com/IceFlameWorm/NLP_Datasets/tree/master/ATEC)
-https://hf.co/datasets/C-MTEB/ATEC
-ATEC NLP sentence pair similarity competition
-s2s
-20,000
-BQ Chen et al. (2018)
-https://hf.co/datasets/C-MTEB/BQ
-Bank Question Semantic Similarity
-s2s
-10,000
-LCQMC Liu et al. (2018)
-https://hf.co/datasets/C-MTEB/LCQMC
-A large-scale Chinese question matching corpus.
-s2s
-12,500
-PAWSX Yang et al. (2019)
-https://hf.co/datasets/C-MTEB/PAWSX
-Translated PAWS evaluation pairs
-s2s
-2,000
-QBQTC Xu et al. (2020a)
-https://hf.co/datasets/C-MTEB/QBQTC
-QQ Browser Query Title Corpus
-s2s
-5,000
-STSB Cer et al. (2017)
-https://hf.co/datasets/C-MTEB/STSB
-Translate STS-B into Chinese
-s2s
-1,360
-STS-22 Muennighoff et al. (2022a)
-https://hf.co/datasets/mteb/sts22-crosslingual-sts
-Chinese news
-p2p
-656
-Appendix B C-MTP Composition
-We mine large-scale pairs of data from various domains. Table 6 shows the details for each data.
-data source
-type of text pairs
-# of pairs
-URL
-cmrc2018
-(query, context)
-9,669
-https://huggingface.co/datasets/cmrc2018
-dureader
-(query, context)
-96,486
-https://github.com/baidu/DuReader
-simclue
-[…]
-388,779
-https://github.com/CLUEbenchmark/SimCLUE
-csl
-(title, abstract)
-394,846
-https://arxiv.org/abs/2209.05034
-amazon_reviews_multi
-(title, body)
-157,762
-https://huggingface.co/datasets/amazon_reviews_multi
-wiki_atomic_edits
-(sentence, edited sentence)
-1,213,688
-https://huggingface.co/datasets/wiki_atomic_edits
-mlqa
-(question, context)
-70,594
-https://huggingface.co/datasets/mlqa
-xlsum
-(title, summary) (title, text)
-89,505
-https://huggingface.co/datasets/csebuetnlp/xlsum
-wudao
-(title, passage)
-37,318,330
-https://data.baai.ac.cn/details/WuDaoCorporaText
-Misc
-60,260,341
-Table 6: Details for each dataset. The Misc data comes from the Internet, including QA, paper, and news data.
-Appendix C English Models
-Using our recipe, we also train a set of English text embedding models presented in Table 7.
-At the time of writing, our English BGE models are state-of-the-art on the English MTEB benchmark (Muennighoff et al., 2022a) across its 56 datasets.
-Our models outperform significantly larger models, such as SGPT Bloom which has 7.1 billion parameters (Muennighoff, 2022; Scao et al., 2022a, b).
-We advance the prior state-of-the-art by an absolute 1.1 (Li et al., 2023b).
-Our training recipe is the same as for our Chinese models, except for the usage of English data.
-We first finetune on unsupervised datasets including datasets like Wikipedia, CC-net, StackExchange, Reddit, S2orc, and datasets from sentence-transformers.
-[^10] https://huggingface.co/datasets/sentence-transformers/embedding-training-data We then further fine-tune on supervised datasets including NLI (Gao et al., 2021c), FEVER (Thorne et al., 2018), NQ (Kwiatkowski et al., 2019), HotpotQA (Yang et al., 2018), Quora, StackExchange Duplicates and MEDI Su et al.
-(2022).
-Model Name Dim. Average Retrieval Cluster Pair CLF Re-rank STS Summarize CLF
-BGE (large) 1024 64.23 54.29 46.08 87.12 60.03 83.11 31.61 75.97
-BGE (base) 768 63.55 53.25 45.77 86.55 58.86 82.4 31.07 75.53
-BGE (small) 384 62.17 51.68 43.82 84.92 58.36 81.59 30.12 74.14
-GTE (large) 1024 63.13 52.22 46.84 85.00 59.13 83.35 31.66 73.33
-GTE (base) 768 62.39 51.14 46.2 84.57 58.61 82.3 31.17 73.01
-E5 (large) 1024 62.25 50.56 44.49 86.03 56.61 82.05 30.19 75.24
-Instructor-XL 768 61.79 49.26 44.74 86.62 57.29 83.06 32.32 61.79
-E5 (base) 768 61.5 50.29 43.80 85.73 55.91 81.05 30.28 73.84
-GTE (small) 384 61.36 49.46 44.89 83.54 57.7 82.07 30.42 72.31
-OpenAI Ada 002 1536 60.99 49.25 45.9 84.89 56.32 80.97 30.8 70.93
-E5 (small) 384 59.93 49.04 39.92 84.67 54.32 80.39 31.16 72.94
-ST5 (XXL) 768 59.51 42.24 43.72 85.06 56.42 82.63 30.08 73.42
-MPNet (base) 768 57.78 43.81 43.69 83.04 59.36 80.28 27.49 65.07
-SGPT Bloom (7.1B) 4096 57.59 48.22 38.93 81.9 55.65 77.74 33.60 66.19
-Table 7: Performance of English Models on MTEB.
+According to the comparison between BGE-pretrain and BGE w.o.
+pre-train, the using of pre-trained text encoder notably improves the retrieval capability, while preserving similar performances on other perspectives.
+5. Conclusion
+In this work, we present C-Pack to advance the progress of general Chinese embedding.
+C-Pack is made up of a comprehensive package of critical resources: 1) The benchmark C-MTEB, covering 6 major tasks of embeddings and 35 datasets, making it the most comprehensive benchmark to evaluate the generality of Chinese text embeddings.
+2) The training data C-MTP, created by the elaborate and comprehensive curation of massive unlabeled corpora and the integration of high-quality labeled data.
+Its unprecedented scale, diversity, and quality contribute to the superior generality of the text embedding models.
+3) The models BGE, which are empirically competitive for both direct usage and downstream fine-tuning.
+The different model sizes provide people with the flexibility to trade off efficiency and embedding quality.
+4) The entire training recipe is also provided along with these resources.
+The public release of C-Pack enjoys a widespread recognition from the technique communities.
+It substantially facilitates the application of general Chinese embedding and also paves the way for its future development.

### v4

@@
 Abstract.
-We introduce C-Pack, a package of resources that significantly advance the field of general text embeddings for Chinese.
+We introduce C-Pack, a package of resources that significantly advances the field of general text embeddings for Chinese.
 C-Pack includes three critical resources.
@@
 Text Embeddings, Training Data, Benchmark, Pre-trained Models
-journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USA.booktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAisbn: 979-8-4007-0431-4/24/07doi: 10.1145/XXXXXX.XXXXXXccs: Information systems Retrieval models and ranking
+journalyear: 2024copyright: acmlicensedconference: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval; July 14–18, 2024; Washington, DC, USAbooktitle: Proceedings of the 47th International ACM SIGIR Conference on Research and Development in Information Retrieval (SIGIR ’24), July 14–18, 2024, Washington, DC, USAdoi: 10.1145/3626772.3657878isbn: 979-8-4007-0431-4/24/07ccs: Information systems Retrieval models and ranking
 1. Introduction
 Text embedding is a long-standing topic in natural language processing and information retrieval.
-By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling (Karpukhin et al., 2020; Lewis et al., 2020; Guu et al., 2020; Xiong et al., 2020).
+By representing texts with latent semantic vectors, text embedding can support various applications, e.g., web search, question answering, and retrieval-augmented language modeling (Zhang et al., 2022; Xiao et al., 2021; Lewis et al., 2020; Xiao et al., 2022a).
 The recent popularity of large language models (LLMs) has made text embeddings even more important.
@@
 We make comprehensive empirical analysis for the value of C-Pack in Section 4.
-Finally, we conclude our work in Section 5.
+Finally, we conclude our work in Section LABEL:sec:conclude.
 Figure 2.
@@
 2. Related Work
-The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering (Karpukhin et al., 2020) but also due to its fundamental role in augmenting large language models (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Shi et al., 2023).
+The importance of general text embedding is widely recognized, not only for its wide usage in typical applications, like web search and question answering (Xiao et al., 2022b; Karpukhin et al., 2020) but also due to its fundamental role in augmenting large language models (Lewis et al., 2020; Guu et al., 2020; Borgeaud et al., 2022; Izacard et al., 2022; Shi et al., 2023).
 Compared with the conventional task-specific methods, the general text embedding needs to be extensively applicable in different scenarios.
@@
 pre-train, the using of pre-trained text encoder notably improves the retrieval capability, while preserving similar performances on other perspectives.
-5. Conclusion
-In this work, we present C-Pack to advance the progress of general Chinese embedding.
-C-Pack is made up of a comprehensive package of critical resources: 1) The benchmark C-MTEB, covering 6 major tasks of embeddings and 35 datasets, making it the most comprehensive benchmark to evaluate the generality of Chinese text embeddings.
-2) The training data C-MTP, created by the elaborate and comprehensive curation of massive unlabeled corpora and the integration of high-quality labeled data.
-Its unprecedented scale, diversity, and quality contribute to the superior generality of the text embedding models.
-3) The models BGE, which are empirically competitive for both direct usage and downstream fine-tuning.
-The different model sizes provide people with the flexibility to trade off efficiency and embedding quality.
-4) The entire training recipe is also provided along with these resources.
-The public release of C-Pack enjoys a widespread recognition from the technique communities.
-It substantially facilitates the application of general Chinese embedding and also paves the way for its future development.